# Dielectric Properties of Single Crystal Calcium Tungstate

### v2

@@ 99 $580987$ $600478$ $2104466$ $2277235$
IV Loss Mechanisms
Loss mechanisms were characterised for the shielded sample, at room-temperature (295K), liquid-helium temperature (4 K), and superfluid helium temperature (100 mK). Fig. 7 displays unloaded $Q$ factors for the two mode families, determined from fitting to the undercoupled [33] Fano resonances of the $S_{21}$ transfer functions. We can infer the effect of loss mechanisms in different frequency regions. At cryogenic temperatures, we see $Q$ factors increase with frequency until they plateau, suggesting that $Q$’s are limited proportionally with cavity wall losses until the WGMs are confined enough to be dominated instead by loss mechanisms within the crystal. The geometry factor quantifies the loss contribution due to the copper cavity walls as
where $\vec{H}$ is the total magnetic field integrated over the volume of the resonator, and $\vec{H}_{t}$ is the tangential magnetic field integrated over the inner surface of the cavity. The geometry factors in units of Ohms ($\Omega$) are tabulated in Table 3. For any select mode, the effect of the cavity wall surface resistance ($R_{s})$ can be calculated with
$Q_{s}=G/R_{s}.$ (8)
where $Q$ is the unloaded Q-factor and $Q_{r}^{-1}$ accounts for the radiative losses, which are negligible in the case of an undercoupled shielded dielectric resonator. Coupling parameters were calculated from the port $1$ and $2$ reflection data and were set at room temperature to be $\mathcal{O}{(10^{-4}-10^{-3})}$ to ensure they remained under-coupled for the full temperature range. $R_{s}$ was determined by measuring $Q$ factors of the resonant modes in the empty oxygen-free cavity at room-temperature. The $R_{s}$ at cryogenic temperature was derived from residual resistivity ratio (RRR) values[34, 35]. At low frequency, the $Q$ factors are dominated by the loss contribution of the $G/R_{s}$ term, as indicated by the little to no change in $Q$ between $4$ K and $100$ mK where $R_{s}$ bottoms out for the case of non-ideal conductors.
The change in trend for frequencies above $6$ GHz, corresponding to $m\geq 6$, indicates dielectric loss prevails in room-temperature and liquid helium temperature conditions. However, at superfluid helium temperature, $Q$s degrade from their $4$ K value, a phenomenon that may be attributed to spins from impurity ions settling into their ground state and becoming the dominant zero-field loss mechanism for resonant frequencies lying in the vicinity of the spin transition energy [36, 37]. Exploration of spin impurities will be conducted in future studies. Therefore, the loss tangent is compared only between $295$ K and $4$ K in the frequency range where dielectric loss dominates.
@@ IV Loss Mechanisms
To calculate the components of the loss tangent for this uniaxial sample, a mode from each fundamental WGM family was selected, and the pair of simultaneous equations was solved, see Eq. 9. The components of the loss tangent of CaWO4 at room-temperature and liquid helium temperature are plotted in Fig. 8. The room-temperature value (295 K) was found to be $\delta_{||}=4.1\pm 1.4\times 10^{-5}, tan$$\delta_{\perp}=3.64\pm 0.92\times 10^{-5}$. At liquid-helium temperature (4 K), the loss tangent is $\delta_{||}=2.05\pm 0.79\times 10^{-7}, tan$$\delta_{\perp}=1.56\pm 0.52\times 10^{-7}$. These values are an order of magnitude worse than reported recently from another sample from a different manufacturer [38]. Results plotted in Fig. 8 indicate a loss peak around 10.5 GHz at 4K, most likely due to an unknown impurity. The effect is broad enough to degrade the Q-factor of both WGE and WGH modes between $9$ to $12$ GHz. This effect highlights the importance of attaining high-purity materials for precision and quantum measurements. Future work will utilise WGM spectroscopy at mK temperature [39] to understand the impurities in this crystal. It has been shown in other crystals that by polarizing all spins at a high magnetic field, the magnetic loss can be made negligible [40].
@@ V Conclusion
We presented a study employing WGM analysis to evaluate the dielectric permittivity of single-crystal CaWO4 from $295$ K to $4$ K and unequivocally determine the relative permittivity tensor at GHz frequencies. These values are important for the future design of microwave circuits that will utilise this crystal. The loss mechanisms associated with the WGM mode analysis and the components of the loss tangent are also reported, and for this sample, were limited by an unknown paramagnetic impurity around 10.5 GHz.

### v3

@@ 99 $580987$ $600478$ $2104466$ $2277235$
The change in trend for frequencies above $6$ GHz, corresponding to $m\geq 6$, indicates dielectric loss prevails in room-temperature and liquid helium temperature conditions. However, at superfluid helium temperature, $Q$ factors degrade from their $4$ K value, a phenomenon that may be attributed to spins from impurity ions settling into their ground state and becoming the dominant zero-field loss mechanism for resonant WGM frequencies lying in the vicinity of the spin transition energy [36, 37]. Exploration of spin impurities will be conducted in future studies. Therefore, the loss tangent is compared only between $295$ K and $4$ K in the frequency range where dielectric loss dominates.
@@
To calculate the components of the loss tangent for this uniaxial sample, a mode from each fundamental WGM family was selected, and the pair of simultaneous equations was solved, see Eq. 9. The components of the loss tangent of CaWO4 at room-temperature and liquid helium temperature are plotted in Fig. 8. The room-temperature value was found to be $\delta^{295\,K}_{||}=4.1\pm 1.4\times 10^{-5}, tan$$\delta^{295\,K}_{\perp}=3.64\pm 0.92\times 10^{-5}$. At liquid-helium temperature the loss tangent was $\delta^{4\,K}_{||}=2.05\pm 0.79\times 10^{-7}, tan$$\delta^{4\,K}_{\perp}=1.56\pm 0.52\times 10^{-7}$. Previous investigations on a different CaWO4 sample indicate that these values are an order of magnitude worse than expected [38]. Results plotted in Fig. 7 suggest a loss peak around 10.5 GHz at 4 K, most likely due to an unknown impurity. The effect is broad enough to degrade the $Q$ factor of both WGE and WGH modes between $9$ to $12$ GHz. This effect highlights the importance of attaining high-purity materials for precision and quantum measurements. Future work will utilise WGM spectroscopy at mK temperature [39] to understand the impurities in this crystal. WGM investigations in other crystals such as SrLaAlO4 demonstrate that by polarising all spins at a high magnetic field, the magnetic loss can be made negligible [40].
@@ V Conclusion
We presented a study employing WGM analysis to evaluate the dielectric permittivity of single-crystal CaWO4 from $295$ K to $4$ K, and determined the relative permittivity tensor at GHz frequencies. These values are important for the future design of microwave circuits that will utilise this crystal. The loss mechanisms associated with the WGM mode analysis and the components of the loss tangent are also reported, and for this sample, were limited by an unknown paramagnetic impurity around 10.5 GHz.